# Complexity Scaling Laws for Neural Models using Combinatorial Optimization

### v1

@@ 1 Introduction
Figure 1: Suboptimality gap, defined as the difference between mean model performance and mean optimal performance, smoothly power decays with respect to model size and compute for both reinforcement learning (RL) and supervised fine-tuning (SFT) in TSP. Fits suggest that SFT is more compute-efficient than RL, and possibly more parameter-efficient as we scale to larger models, with faster decay toward optimal performance (larger $\alpha$). Top left: Suboptimality w.r.t. parameters ($N$) for models evaluated near convergence. Right: Suboptimality w.r.t. compute ($C$) evaluated throughout training, where the compute-efficient frontier power decays. Note that the compute axis for SFT has been stretched for easier viewing. Bottom left: Optimal model size follows power growth w.r.t. compute budget. This relationship is strikingly consistent between domains [13, 4, 14].
@@ 2.1 Data design
All experiments were performed using the Symmetric Euclidean Traveling Salesman Problem. One problem instance consists of $n$ $\mathbb{R}^{d}$ coordinates uniformly sampled between $[0,1]$ for each dimension $d$ (illustrated in Figure 2). TSP’s $O(n!)$ combinatorial solution space is defined as all closed loops that visit each node exactly once. Such loops are called tours, and the objective is to find the minimum-length tour. In combinatorial optimization language, tour length is referred to as cost (or fitness when maximizing), and the cost surface w.r.t. the solution space is referred to as the cost landscape.
@@ 2.5 Scaling experiments
Problem complexity scaling: For TSP node and spatial dimension scaling laws, we fix model width thereby fixing model capacity as we scale problem complexity. For node scaling, we train one RL and one SFT model per node scale. We study ten scales, $n\in\{5,10,15,...,50\}$. Symmetric TSP has $\nicefrac{{1}}{{2}}\,(n-1)!$ possible tours, [^3] Starting node insensitivity subtracts 1, and tour reversal insensitivity divides by 2. so these scales range from 12 solutions up to approximately $3\times 10^{62}$. For spatial dimension scaling, we train one RL model per scale and forgo SFT training given the scarcity of optimal data as detailed in Section 2.2. We study $d\in\{2,3,...,12,15,20,30,40,50,100\}$ over two node scales $n\in\{10,20\}$, which allows us to better interpret otherwise ambiguous trends.
@@ 3 Neural scaling laws for combinatorial optimization
A broader implication of Figure 1 is that TSP tour length is a natural performance metric [14]. For TSP, RL obtains smooth scaling laws directly for the return signal with high precision. For more complex environments with sequential dynamics, return-based RL scaling laws significantly lose predictive power [22, 9, 23, 24].
@@ 4.1 Solution space scaling
Even if we assume that $\alpha$ estimates are imprecise, the growth of suboptimality is clearly superlinear. This observation presents a subtle contradiction. Figure 4 shows the suboptimality of random tour length grows almost linearly with node scale. Indeed, this growth is provably $O(n)$ as $n\to\infty$. [^5] Lemma 1 shows that expected random tour length grows linearly w.r.t. $n$, while the Beardwood-Halton-Hammersley Theorem [25] proves that optimal tour length is asymptotically proportional to $\sqrt{n}$ as $n\to\infty$. Thus, the obtained scaling laws imply model performance will eventually be worse than random performance, and with enough scale, exceed even the maximum possible tour length (also $O(n)$). Instead, we expect this superlinearity to break down before exceeding random performance, and we expect the obtained scaling laws to become pessimistic estimates at large node scales.
@@
Figure 4: Node and spatial dimension scaling have distinct effects on the achievable performance span, the suboptimality gap of random performance. Left: Mean optimal tour length closely follows sublinear power growth w.r.t. either problem scale. Mean random tour length grows linearly w.r.t. nodes, and sublinearly w.r.t. dimensions at rate similar to optimal tour length growth. Each sublinear trend approaches square root growth in the limit ($\alpha=0.5$; proof in Appendix I). Right: Suboptimality of random performance w.r.t. nodes is polynomial but approximately linear being dominated by random tour length growth. Suboptimality of random performance w.r.t. dimensions produces a small, transient increase then decrease, but is provably constant in the limit (Theorem 6).
@@ 4.2 Representation space scaling
The right plot in Figure 3 shows suboptimality for RL models near convergence as we scale the number of TSP spatial dimensions. Both 10-node and 20-node experiments produce smooth bounded growth described by negative exponential decay toward a fitted asymptote $\beta$. We found that almost any curve can fit the trend of the 10-node experiment since its growth is closer to being linear, but the 20-node experiment is more discriminative. Unbounded power growth produces a markedly poor fit given the observed trend is visibly convergent. Power decay regression obtains smaller residuals but appears to converge slightly slower than the observed growth. Further, power decay fits produce a likely contradiction: the bound obtained for the 10-node experiment ($\beta_{\alpha}\approx 0.31$) is higher than that obtained for the more complex 20-node counterpart ($\beta_{\alpha}\approx 0.25$). This implies that a lower suboptimality is achieved in the limit as $d\to\infty$ by raising the baseline solution complexity, unless increased node scale fundamentally alters the form of scaling w.r.t. dimensions. In contrast, we find that exponential decay is most predictive [^6] We also tested decays with the quasi-polynomial form $d^{-\log_{\psi}d}$ and the subexponential form $\psi^{{-d}^{\phi}}$, $\phi\in(0,1]$ but found extreme log base fits for the former and no visual improvement from either form. while maintaining sound asymptote ordering between 10- and 20-node experiments (the $\beta_{\psi}$ values shown). Like $\alpha$ in node scaling fits, $\beta_{\psi}$ is sensitive to the suboptimality values obtained at larger scales. Because higher-dimensional runs require more training to converge, $\beta_{\psi}$ estimates are likely imprecise, so the exact values shown may not be meaningful.
@@ 5 Interpretable complexity scaling with local search
We now reverse engineer the complexity scaling laws introduced in Section 4 using local search, studying comparable trends produced with a white-box algorithm. These similarities do not directly imply that model inference and local search share underlying mechanisms. Instead, they provide potential topics to explore in future work, which we identify throughout this section. We evaluate local search using the same datasets discussed in Section 2.4. For each problem, we initialize search with a random tour and then generate a locally optimum tour with the simple, well-studied 2-opt search move [31], performing gradient descent through the cost landscape. The landscape properties discussed at a high level below are detailed in Appendix D.
@@ 6 Related work
Neural scaling laws where cross-entropy loss smoothly improves with model size, dataset size, or compute budget were popularized by Kaplan et al.’s study of large language models [3], which also first demonstrated the separable nature of embedding parameters. Henighan et al. [4] generalize neural scaling laws to other autoregressive generative modeling domains, showing that optimal model size w.r.t. compute budget is remarkably consistent. Supervised learning constitutes the majority of recent research on neural scaling laws [21, 33, 34, 5, 6, 35, 36]. Several theories have been proposed to explain the relationships between loss and parameters [37], between loss and data [38], and for both in conjunction [39, 40]. Most recently, Bordelon et al.’s dynamical mean field theory [41] recovers phenomena from parameter, data, and compute scaling. The Quantization Hypothesis [40, 42] also explains emergent abilities [43] by positing that there is an underlying discreteness to all abilities that deep models learn. However, Schaeffer et al. [44] show that emergent abilities often reflect the chosen performance metric rather than deep learning itself.
@@ 7 Discussion
Infinite-compute performance can only be estimated. With modest resources, we could not experiment on large models and problem scales. The strongest limitation of this work is its ability to extrapolate, especially for node scaling trends which must break down. We are also limited in precision when comparing SFT and RL scaling law fits given that fewer SFT models converge to trend with one epoch of training, and we train only a single seed per scale. Another limitation arises from Euclidean TSP being easier than most NP-hard problems, having polynomial-time approximation schemes (PTAS) with respect to the number of nodes [52]. This property of TSP should be considered when hypothesizing how these results generalize to other combinatorial optimization problems (and beyond).
@@
Despite TSP’s relative simplicity, there is some basis to be optimistic for parameter-constrained complexity scaling laws in other domains. The TSP cost landscape was once widely accepted to be approximately globally convex [53, 54], but more recent work has found this convexity to be increasingly coarse [55, 56, 57]. Even so, in expectation, smooth complexity scaling laws emerge. This macroscopic order despite microscopic disorder mirrors the nature of neural scaling laws and supports the hypothesis that complexity scaling laws can be obtained for less structured tasks. Nevertheless, solution complexity and representation complexity are often tightly coupled multivariate functions in domains like language modeling, vision, robotics, and so on. For instance, in language modeling, vocabulary size affects both the number of token representations and the space of possible texts to generate. Identifying scalable yet precise measures of problem complexity is one challenge for future work in these domains. Another challenge arises for tasks where relevant performance metrics do not smoothly improve, as is often the case for sequential decision processes and tasks with discontinuous performance metrics. Though the latter has recently been addressed for language modeling with Token Edit Distance [44], an edit-space analogue to suboptimality.
@@
Without an underlying theory that explains parameter-constrained complexity scaling laws, interpreting and relating findings between domains will be more difficult. Explaining each form of scaling is the most obvious gap. Extending Section 5 to reverse engineer scaling laws with a white-box machine learning model is one possible approach. We are especially interested in whether model size predicts suboptimality growth rates like $\alpha$ and $\psi$, or $\beta$ limits for spatial dimension scaling, and explaining the underlying nature if so. Another gap in our current understanding is what determines an embedding bottleneck. Widening the coordinate projection layer forever is probably insufficient to avoid an embedding bottleneck while dimension scaling. Extending the positional encoding sequence forever may be insufficient to avoid one while node scaling.
Compute-limited complexity scaling:
While this paper focuses on bottleneck from model capacity, we could have instead chosen to limit compute budget or dataset size (with early stopping) and observe the patterns that emerge with increased problem complexity, if any. For example, Jones [10] shows that we can predict model performance as a bivariate function of train-time compute and test-time compute for a single Hex board size. If this neural scaling law can be generalized over problem scale, that would have many practical use cases.
@@ Appendix A Scaling law fit constants
RL and SFT scaling laws may have non-trivial imprecision due to training with a single seed per scale. This should be considered when extrapolating, or when comparing values between RL and SFT such as $\alpha$ in node and parameter scaling fits.
@@
2-opt $20$ $0.254$ $98.6$ $0.251$ $38.0$

### v2

@@ 1 Introduction
Figure 1: Suboptimality gap, defined as the difference between mean model performance and mean optimal performance, smoothly power decays with respect to model size and compute for both reinforcement learning (RL) and supervised fine-tuning (SFT) in TSP. Fits suggest that SFT is more compute-efficient than RL, and possibly more parameter-efficient as we scale to larger models, with faster decay toward optimal performance (larger $\alpha$). Top left: Suboptimality w.r.t. parameters (NN) for models evaluated near convergence. Right: Suboptimality w.r.t. compute ($C$) evaluated throughout training, where the compute-efficient frontier power decays. Note that the compute axis for SFT has been stretched for easier viewing. Bottom left: Optimal model size follows power growth w.r.t. compute budget. This relationship is strikingly consistent between domains [13, 4, 14].
@@ 2.1 Data design
All experiments were performed using the Symmetric Euclidean Traveling Salesman Problem. One problem instance consists of $n$ $\mathbb{R}^{d}$ coordinates uniformly sampled between $[0,1]$ for each of the $d$ dimensions (illustrated in Figure 2). TSP’s $O(n!)$ combinatorial solution space is defined as all closed loops that visit each node exactly once. Such loops are called tours, and the objective is to find the minimum-length tour. In combinatorial optimization language, tour length is referred to as cost (or fitness when maximizing), and the cost surface w.r.t. the solution space is referred to as the cost landscape.
@@ 2.5 Scaling experiments
Problem complexity scaling: For TSP node and spatial dimension scaling laws, we fix model width thereby fixing model capacity as we scale problem complexity. For node scaling, we train one RL and one SFT model per node scale. We study ten scales, $n\in\{5,10,15,...,50\}$. Symmetric TSP has $\nicefrac{{1}}{{2}}\,(n-1)!$ possible tours [^3] From $n!$ permutations, starting node insensitivity yields $n$ equivalent tours per permutation and tour reversal insensitivity yields $2$ equivalent tours per permutation (dividing by $n$ and $2$, respectively). for $n\geq 3$, so these scales range from 12 solutions up to approximately $3\times 10^{62}$. For spatial dimension scaling, we train one RL model per scale and forgo SFT training given the scarcity of optimal data as detailed in Section 2.2. We study $d\in\{2,3,...,12,15,20,30,40,50,100\}$ over two node scales $n\in\{10,20\}$, which allows us to better interpret otherwise ambiguous trends.
@@ 3 Neural scaling laws for combinatorial optimization
A broader implication of Figure 1 is that TSP tour length is a natural performance metric [14]. For TSP, RL obtains smooth scaling laws directly for the return signal with high precision. For more complex environments with sequential dynamics, return-based RL scaling laws have considerable variance and lose predictive power [22, 9, 23, 24].
@@ 4.1 Solution space scaling
Even if we assume that $\alpha$ estimates are imprecise, the growth of suboptimality is clearly superlinear. This observation presents a subtle contradiction. Figure 4 shows that the suboptimality of random tour length grows almost linearly with node scale. Indeed, this growth is provably $O(n)$ as $n\to\infty$. [^5] Lemma 1 shows that expected random tour length grows linearly w.r.t. $n$, while the Beardwood-Halton-Hammersley Theorem [25] proves that optimal tour length is asymptotically proportional to $\sqrt{n}$ as $n\to\infty$. Thus, the obtained scaling laws imply model performance will eventually be worse than random performance, and with enough scale, exceed even the maximum possible tour length (also $O(n)$). Instead, we expect this superlinearity to break down before exceeding random performance, and we expect the obtained scaling laws to become pessimistic estimates at large node scales.
@@
Figure 4: Node and spatial dimension scaling have distinct effects on the achievable performance span, the suboptimality gap of random performance. Left: Mean optimal tour length closely follows sublinear power growth w.r.t. either problem scale. Mean random tour length grows linearly w.r.t. nodes, and sublinearly w.r.t. dimensions at rate similar to optimal tour length growth. Each sublinear trend approaches square root growth in the limit ($\alpha=0.5$; proof in Appendix I). Right: Suboptimality of random performance w.r.t. nodes is polynomial but approximately linear, being dominated by random tour length growth. Suboptimality of random performance w.r.t. dimensions produces a small, transient increase then decrease, but is provably constant in the limit (Theorem 6).
@@ 4.2 Representation space scaling
The right plot in Figure 3 shows suboptimality for RL models near convergence as we scale the number of TSP spatial dimensions. Both 10-node and 20-node experiments produce smooth bounded growth described by negative exponential decay toward a fitted asymptote $\beta$. We found that almost any regression model results in a good fit of the 10-node experiment since its growth is closer to being linear, but the 20-node experiment is more discriminative. Unbounded power growth produces a markedly poor fit given the observed trend is visibly convergent. Power decay regression obtains smaller residuals but appears to converge slightly slower than the observed growth. Further, power decay fits produce a likely contradiction: the bound obtained for the 10-node experiment ($\beta_{\alpha}\approx 0.31$) is higher than that obtained for the more complex 20-node counterpart ($\beta_{\alpha}\approx 0.25$). This implies that a lower suboptimality is achieved in the limit as $d\to\infty$ by raising the baseline solution complexity, unless increased node scale fundamentally alters the form of scaling w.r.t. dimensions. In contrast, we find that exponential decay is most predictive [^6] We also tested decays with the quasi-polynomial form $d^{-\log_{\psi}d}$ and the subexponential form $\psi^{{-d}^{\phi}}$, $\phi\in(0,1]$ but found extreme log base fits for the former and no visual improvement from either form. while maintaining sound asymptote ordering between 10- and 20-node experiments (the $\beta_{\psi}$ values shown). Like $\alpha$ in node scaling fits, $\beta_{\psi}$ is sensitive to the suboptimality values obtained at larger scales. Because higher-dimensional runs require more training to converge, $\beta_{\psi}$ estimates are likely imprecise, so the exact values shown may not be meaningful.
@@ 5 Interpretable complexity scaling with local search
We now reverse engineer the complexity scaling laws introduced in Section 4 using local search, studying comparable trends produced with a white-box algorithm. These similarities do not directly imply that model inference and local search share underlying mechanisms. Instead, they provide potential topics to explore toward the development of a formal theory for parameter-constrained deep learning in future work. We evaluate local search using the same datasets discussed in Section 2.4. For each problem, we initialize search with a random tour and then generate a locally optimum tour with the simple, well-studied 2-opt search move [31], performing gradient descent through the cost landscape. The landscape properties discussed at a high level below are detailed in Appendix D.
@@ 6 Related work
Neural scaling laws where cross-entropy loss smoothly improves with model size, dataset size, or compute budget were popularized by Kaplan et al.’s study of large language models [3], which also first demonstrated the separable nature of embedding parameters. Henighan et al. [4] generalize neural scaling laws to other autoregressive generative modeling domains, showing that optimal model size w.r.t. compute budget is remarkably consistent. Sorscher et al. [33] demonstrate that an exponential decay of test error w.r.t. data can be achieved using data pruning (instead of the typical power law scaling [3, 5, 6]). Analogously, Frantar et al. [34] show that network sparsity influences scaling behavior. Supervised learning constitutes the majority of recent research on neural scaling laws [3, 4, 21, 33, 34, 35, 5, 6, 36, 37].
@@
Neural scaling law theory:
Several theories have been proposed to explain the relationships between loss and parameters [41], between loss and data [42], and for both in conjunction [43, 44]. Most recently, Bordelon et al.’s dynamical mean field theory [45] recovers phenomena from parameter, data, and compute scaling. The Quantization Hypothesis [44, 38] also explains emergent abilities [46] by positing that there is an underlying discreteness to all abilities that deep models learn. However, Schaeffer et al. [47] show that emergent abilities often reflect the chosen performance metric rather than deep learning itself.
@@ 7 Discussion
Infinite-compute performance can only be estimated. With modest resources, we could not experiment on larger models and problem scales. This is especially important to consider when extrapolating node scaling trends, since they must eventually break down (possibly well before the contradiction detailed in Section 4.1). We are also limited in precision when comparing SFT and RL scaling law fits given that fewer SFT models converge to trend with one epoch of training, and we train only a single seed per scale. [^7] We provide leave-one-out fit statistics in Appendix A. Nevertheless, the main limitation of this work is its focus on a single problem. Euclidean TSP is easier than most NP-hard problems, in part due to its cost metric satisfying the triangle inequality [49]. Euclidean TSP also admits polynomial-time approximation schemes (PTAS), allowing near-optimal tours to be computed in polynomial time with respect to the number of nodes [53]. These properties should be considered when hypothesizing how our findings generalize to other combinatorial optimization problems and beyond. Future work would need to demonstrate compelling generalization before complexity scaling laws can be substantiated as a general principle.
@@
Despite TSP’s relative simplicity, there is some basis to be optimistic for parameter-constrained complexity scaling laws in other tasks. The TSP cost landscape was once widely accepted to be approximately globally convex [54, 55], but more recent work has found this convexity to be increasingly coarse [56, 57, 58]. Even so, in expectation, smooth complexity scaling laws emerge. This macroscopic order despite microscopic disorder mirrors the nature of neural scaling laws and supports the hypothesis that complexity scaling laws can be obtained for less structured tasks. A natural next step would be studying parameter-constrained complexity scaling with more elaborate Euclidean combinatorial optimization problems where the solution space and the representation space remain independently scalable. Potential testbeds include other routing problems like the Vehicle Routing Problem (VRP) [59], the Orienteering Problem (OP) [60], and the Ring Star Problem (RSP) [61]. Given the similarity between the RL node scaling curve and its SFT counterpart (Figure 3 left), we especially encourage future study of algorithm insensitivity for parameter-constrained complexity scaling in other problems.
Real-world domains:
Identifying scalable yet precise measures of problem complexity is very challenging in domains like language modeling, vision, robotics, and so on. Solution complexity and representation complexity are often tightly coupled. For instance, in language modeling, vocabulary size affects both the number of token representations and the space of possible texts to generate. Discriminative NLP tasks like multiple choice question answering maintain separability of vocabulary size and number of classes, but at the expense of diminished relevance to generative modeling. Another challenge arises for real-world tasks where relevant performance metrics do not smoothly improve, as is often the case for sequential decision processes and tasks with discontinuous performance metrics. However, the latter has recently been addressed for language modeling with Token Edit Distance [47], an edit-space analogue to suboptimality.
@@
Without an underlying theory that explains parameter-constrained complexity scaling laws, we can only obtain a limited understanding of the conditions in which they apply and the types of problems they generalize to. Relating findings between domains will require significant speculation and extrapolations may be unreliable. Explaining each form of scaling is the most obvious gap. We are especially interested in whether model size predicts suboptimality growth rates like $\alpha$ and $\psi$, or $\beta$ limits for spatial dimension scaling, and explaining the underlying nature if so. Extending Section 5 to reverse engineer scaling laws with a white-box machine learning model is one possible approach. Identifying and explaining counter-examples is another potentially useful approach. For example, if network sparsity were to influence parameter-constrained complexity scaling laws (analogous to how sparsity influences neural scaling laws [34]), a theory explaining this distinction may also address related open questions. Separately, another gap in our current understanding is what determines an embedding bottleneck. Widening the coordinate projection layer forever is probably insufficient to avoid an embedding bottleneck while dimension scaling. Extending the positional encoding sequence forever may be insufficient to avoid one while node scaling.
Practical implications:
Deep methods that excel at solving TSP usually excel at solving similar routing problems like VRP and OP [62, 63, 64], both of which have well-studied practical variants [59, 60]. Practical use cases for parameter-constrained complexity scaling laws are more challenging to identify given that training until convergence is usually compute-inefficient. One use case would be predicting performance for applications mainly constrained by model size, such as edge applications requiring fast inference with limited hardware acceleration. Another potential use case is algorithm benchmarking, where we can estimate the limit of performance for several algorithms across a range of problem complexities without training models on larger problem scales. Such benchmarks can also provide useful insights like sample efficiency tradeoffs with respect to problem complexity. Our single-seed node scaling results for RL and SFT demonstrate that SFT reaches intermediate suboptimality values faster at the scales evaluated (Figure 8 in Appendix B). Future work could use multiple-seed training to quantify this advantage with high statistical confidence.
Finite-compute complexity scaling:
While this paper focuses on bottleneck from model capacity, we could have instead chosen to limit compute budget or dataset size (with early stopping) and observe the patterns that emerge with increased problem complexity, if any. Compute-constrained and data-constrained complexity scaling laws would have many practical use cases. For example, Jones [10] shows that we can predict model performance as a bivariate function of train-time compute and test-time compute for a single Hex board size. If this neural scaling law can be generalized over problem scale, one could estimate for complex tasks the tradeoff between train-time and test-time compute using only cheap experiments.
@@ Appendix A Scaling law fit constants
RL and SFT scaling laws may have non-trivial imprecision due to training with a single seed per scale. This should be considered when extrapolating, or when comparing values between RL and SFT such as $\alpha$ in node scaling and optimal model size fits. To mitigate this concern, we performed jackknife (leave-one-out) resampling of RL and SFT fits and provide growth rate statistics in Table 4. We provide individual leave-one-out (LOO) fits in the project repository. While LOO distributions are distinct from those obtained using multiple seeds, they quantify sensitivity to individual training runs.
Optimal model size fits are relatively sensitive due to the frontier overlap between adjacent model sizes. We observe higher variance and the largest SFT LOO $\alpha$ exceeds the smallest RL LOO $\alpha$ (despite the full-data RL $\alpha$ being larger than the full-data SFT $\alpha$). Node-scaling fits are relatively sensitive to the suboptimalities obtained at the largest node scales. For RL, we obtain $\alpha$ values as large as $1.95$ when fixing the $\gamma$ offset at its full-data value and $2.22$ when fitting $\gamma$. Demonstrated by the large positive skewness, these larger $\alpha$ fits occur at larger LOO node scales. However, fitting $\gamma$ can shift the offset more than $3$ nodes and we know near optimality is achieved at $5$ nodes, so fixed-$\gamma$ LOO variance is likely closer to multi-seed variance. Remaining LOO fits are quite stable, and we found that $\beta$ asymptote fits for dimension scaling experiments are nearly constant (roughly $\pm 0.001$).
@@
2-opt 2020 $0.254$ $98.6$ $0.251$ $38.0$
Table 4: Leave-one-out (LOO) statistics for RL and SFT fits of exponent $\alpha$ and base $\psi$.
LOO indices (idx) are in ascending scale order.
Scaling law Algorithm Fit $\mu$ $\sigma$ Skew Min (idx) Max (idx)
Parameters ($N$) RL $\alpha$ $0.579$ $0.013$ 1.9-1.9 0.545(0)0.545\,(0) 0.593(4)0.593\,(4)
SFT $\alpha$ $0.718$ $0.015$ $0.87$ 0.701(2)0.701\,(2) 0.745(0)0.745\,(0)
Compute ($C_{\min}$) RL $\alpha$ $0.440$ $0.011$ $0.83$ 0.423(5)0.423\,(5) 0.463(0)0.463\,(0)
SFT $\alpha$ $0.554$ $0.011$ 0.28-0.28 0.536(2)0.536\,(2) 0.570(0)0.570\,(0)
Optimal model size RL $\alpha$ $0.811$ $0.064$ 0.47-0.47 0.690(7)0.690\,(7) 0.903(5)0.903\,(5)
SFT $\alpha$ $0.655$ $0.071$ 0.22-0.22 0.556(4)0.556\,(4) 0.734(2)0.734\,(2)
Nodes ($n$) - fix $\gamma$ RL $\alpha$ $1.87$ $0.032$ $1.4$ 1.83(6)1.83\,(6) 1.95(8)1.95\,(8)
SFT $\alpha$ $1.71$ $0.059$ $1.4$ 1.65(3)1.65\,(3) 1.83(5)1.83\,(5)
Nodes ($n$) - fit $\gamma$ RL $\alpha$ $1.90$ $0.116$ $2.2$ 1.81(0)1.81\,(0) 2.22(8)2.22\,(8)
SFT $\alpha$ $1.69$ $0.134$ $0.46$ 1.48(0)1.48\,(0) 1.94(5)1.94\,(5)
Dimensions ($d$)
10 $n$
20 $n$
RL $\psi$ $1.18$ $0.007$ 1.0-1.0 1.16(0)1.16\,(0) 1.19(5)1.19\,(5)
RL $\psi$ $1.68$ $0.023$ 0.46-0.46 1.62(2)1.62\,(2) 1.71(0)1.71\,(0)